### v1

@@ Appendix A Proofs of Main Theorem
¯$\overline{\mathbb{P}}\left(\left\|\sum_{t=1}^{T}\left[\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}-\overline{\mathbb{E}}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right]\right\|_{2}\geq x\right)\leq 2d\exp\left(-\frac{x^{2}/2}{\sigma_{1}^{2}+N_{1}x/3}\right).$
Choosing $x\asymp\sqrt{d}\log(d)\|\sum_{i=1}^{r}f_{it}a_{i}\|_{2}+\sqrt{\log(d)}\sigma_{2}$, with probability at least $1-d^{-c_{4}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}-\overline{\mathbb{E}}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2}\leq C_{4}\frac{\sqrt{\log(d)}}{T}\sigma_{2}+C_{4}\frac{\sqrt{d}\log(d)\|\sum_{i=1}^{r}f_{it}a_{i}\|_{2}}{T}.$
As $\sqrt{r}\log(T)^{1/\gamma_{1}}\lesssim\sqrt{d}$, by Assumption 4.2, with probability at least $1-T^{-c_{5}}$,
$\left\|\sum_{i=1}^{r}f_{it}a_{i}\right\|_{2}=\|F_{t}^{\top}A^{\top}\|_{2}\leq(1+\delta)\|F_{t}\|_{2}\lesssim\sqrt{r}(\log(T))^{1/\gamma_{1}}\sqrt{\lambda_{1}}\lesssim\sqrt{d\lambda_{1}}.$
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}-\mathbb{E}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2}\leq C_{5}\sqrt{\frac{d\log(d)}{T}}\sqrt{\lambda_{1}}+C_{5}\cdot\frac{d\log(d)\sqrt{\lambda_{1}}}{T}.$
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}\geq x\right)\leq\mathbb{P}(|M|>C)+\mathbb{P}\left(|M|<C,|M|\max_{t}\left\|\sum_{i=1}^{r}f_{it}a_{i}\right\|_{2}\cdot\|e_{t}\|_{2}\geq x\right).$
Choosing $x\asymp d\sqrt{\lambda_{1}}$, we have with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}-T^{-c_{5}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}\leq C_{6}\cdot\frac{d\sqrt{\lambda_{1}}}{T}$
@@
$\|\Delta_{3}\|_{2}\leq$ $\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}-\mathbb{E}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}\}}\right\|_{2}$
$+\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}+\left\|\mathbb{E}\sum_{i=1}^{r}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}\}}\right\|_{2}$
$\leq$ $C_{7}\sqrt{\frac{d\log(d)}{T}}\sqrt{\lambda_{1}}+C_{7}\cdot\frac{d\log(d)\sqrt{\lambda_{1}}}{T}.$
@@
Let $\lambda_{1}/\lambda_{r}=w$. Assume $d_{1}\lesssim\lambda_{1}$ and $\delta_{1}^{2}w\leq c$ for a sufficiently small positive constant $c$. Apply random projection in Procedure 2 to the whole sample covariance tensor $\widehat{\Sigma}$ with $L\geq Cd^{2}\vee Cdr^{2w^{2}}$. Denote the estimated CP basis vectors as $\widetilde{a}_{\ell k}$, for $1\leq\ell\leq L,1\leq k\leq K$. Then in an event with probability at least $1-T^{-c}-d^{-c}$, we have for any CP factor loading vectors tuple $(a_{ik},1\leq k\leq K)$, there exist $j_{i}\in[L]$ such that
@@
$\left\|\widetilde{a}_{j_{i},1}\widetilde{a}_{j_{i},1}^{\top}-a_{i1}a_{i1}^{\top}\right\|_{2}$ $\leq\psi_{i}+(\delta/\delta_{1})w+\psi_{i}^{K-1}w,$ (47)
where $\psi_{i}=CR^{(0)}/\Theta_{ii}+\delta_{1}^{2}w$, $R^{(0)}=\phi^{(0)}$ is defined in (19) in Theorem 4.1, and $1\leq i\leq r$.
@@
Without loss of generality, assume $\Theta_{11}\geq\Theta_{22}\geq\cdots\geq\Theta_{rr}$. Then $\Theta_{rr}\geq\lambda_{r},\Theta_{11}\leq\lambda_{1}$. Let $\widehat{\Theta}_{ij}=T^{-1}\sum_{t=1}^{T}f_{it}f_{jt}$ and $\Sigma_{{\cal E}}=\mathbb{E}{\cal E}_{t}\otimes{\cal E}_{t}$. Write
@@
$\quad+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{it}\otimes_{k=1}^{K}a_{ik}\otimes{\cal E}_{t}+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{it}{\cal E}_{t}\otimes_{k=K+1}^{2K}a_{ik}+\left(\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes{\cal E}_{t}-\Sigma_{{\cal E}}\right)+\Sigma_{{\cal E}}$
@@
$\leq C\sqrt{d_{1}}\left(\sqrt{\frac{d\log(d)}{T}}+\frac{d\log(d)}{T}\right)\cdot\sqrt{\log(d)},$
@@
$\left\|(\widetilde{\Delta}_{1}+\widetilde{\Delta}_{2}+\widetilde{\Delta}_{3}+\widetilde{\Delta}_{4})\times_{3}\times_{4}\theta\right\|_{2}\leq$ $C\lambda_{1}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right)\cdot\sqrt{\log(d)}$
$+C\lambda_{1}^{1/2}\left(\sqrt{\frac{d\log(d)}{T}}+\frac{d\log(d)}{T}\right)\cdot\sqrt{\log(d)}.$
@@
$\geq(1-2\delta_{1}^{2}w)\left(\sqrt{2\log(L)}-\frac{\log\log(L)}{4\sqrt{\log(L)}}-\sqrt{2\log(8)}\right)\Theta_{ii}-\left(\sqrt{4\log(r)}+\sqrt{2\log(8)}\right)w\Theta_{ii}$
@@
with probability at least $\frac{1}{2}$, by letting $L\geq Cd\vee Cr^{2w^{2}}$.
Since $\theta_{\ell}$ are independent samples, we instead take $L_{i}=L_{i1}+\cdots+L_{iM}$ for $M=\lceil C_{1}\log(d)/\log(2)\rceil$ and $L_{i1},...,L_{iM}\geq Cd\vee Cr^{2w^{2}}$. We define
@@
$\left\|\widetilde{a}_{\ell_{*},k}\widetilde{a}_{\ell_{*},k}^{\top}-a_{ik}a_{ik}^{\top}\right\|_{2}\leq\frac{CR^{(0)}}{\Theta_{ii}}+\delta_{1}^{2}w,\qquad 2\leq k\leq K.$ (52)
Now consider to obtain $\widetilde{a}_{\ell_{*},1}$. Write $\psi_{i}=CR^{(0)}/\Theta_{ii}+\delta_{1}^{2}w$. Note that
@@
$\leq C_{K}(\delta/\delta_{1}+\psi_{i}^{K-1})w\Theta_{ii}.$
@@
$\left\|\widetilde{a}_{\ell_{*},1}\widetilde{a}_{\ell_{*},1}^{\top}-a_{i1}a_{i1}^{\top}\right\|_{2}\leq\frac{CR^{(0)}}{\Theta_{ii}}+(\delta/\delta_{1})w+\psi_{i}^{K-1}w.$ (54)
Repeat the same argument again for all $1\leq i\leq r$ factors, and let $L=\sum_{i}L_{i}\geq Cd^{2}\vee Cdr^{2w^{2}}\geq Cdr\log(d)\vee Cr^{2w^{2}+1}\log(d)$. We have, in the event $\Omega_{0}\cap\Omega_{1}\cap\cdots\cap\Omega_{r}$ with probability at least $1-T^{-c}-d^{-c}$, (52) and (54) hold for all $i$.
@@
Recall $\widehat{A}_{k}^{(m)}=(\widehat{a}_{1k}^{(m)},\ldots,\widehat{a}_{rk}^{(m)})\in\mathbb{R}^{d_{k}\times r}$, $\widehat{\Sigma}_{k}^{(m)}=\widehat{A}_{k}^{(m)\top}\widehat{A}_{k}^{(m)}$, and $\widehat{B}_{k}^{(m)}=\widehat{A}_{k}^{(m)}(\widehat{\Sigma}_{k}^{(m)})^{-1}=(\widehat{b}_{1k}^{(m)},...,\widehat{b}_{rk}^{(m)})\in\mathbb{R}^{d_{k}\times r}$. Let $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|F_{t},1\leq t\leq T)$ and ¯$\overline{\mathbb{P}}(\cdot)=\mathbb{P}(\cdot|F_{t},1\leq t\leq T)$. Also let
$\overline{\lambda}_{i}=\frac{1}{T}\sum_{t=1}^{T}f_{it}f_{it}.$
@@
$=\sum_{i=1}^{r}\overline{\lambda}_{i}\otimes_{k=1}^{2K}a_{ik}+\sum_{i\neq j}^{r}\frac{1}{T}\sum_{t=1}^{T}f_{it}f_{jt}\otimes_{k=1}^{K}a_{ik}\otimes_{k=K+1}^{2K}a_{jk}+\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes{\cal E}_{t}$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{it}\otimes_{k=1}^{K}a_{ik}\otimes{\cal E}_{t}+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}f_{it}{\cal E}_{t}\otimes_{k=K+1}^{2K}a_{ik}$
@@
Let $g_{i\ell}=b_{i\ell}/\|b_{i\ell}\|_{2}$ and $\widehat{g}_{i\ell}^{(m)}=\widehat{b}_{i\ell}^{(m)}/\|\widehat{b}_{i\ell}^{(m)}\|_{2}$. Given $\widehat{a}_{i\ell}^{(m)}$ ($1\leq i\leq r,1\leq\ell\leq K$), the $(m+1)$th iteration produces estimates $\widehat{a}_{ik}^{(m+1)}$, which is the top left singular vector of $\widehat{\Sigma}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{b}_{i\ell}^{(m)\top}$, or equivalently $\widehat{\Sigma}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top}$. Note that $\widehat{\Sigma}=\sum_{j=1}^{r}\overline{\lambda}_{j}\otimes_{\ell=1}^{2K}a_{j\ell}+\Psi$, with $a_{j,\ell+K}=a_{j\ell}$. The “noiseless" version of this update is given by
$\widehat{\Sigma}\times_{\ell\in[2K]\backslash\{k,K+k\}}g_{i\ell}^{\top}=\overline{\lambda}_{i}a_{ik}a_{ik}^{\top}+\Psi\times_{\ell\in[2K]\backslash\{k,K+k\}}g_{i\ell}^{\top}.$ (57)
@@
$\widehat{\Sigma}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top}=\sum_{j=1}^{r}\widetilde{\lambda}_{j,i}a_{jk}a_{jk}^{\top}+\Psi\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top},$
@@
$\widetilde{\lambda}_{j,i}=\overline{\lambda}_{j}\prod_{\ell\in[2K]\backslash\{k,K+k\}}a_{j\ell}^{\top}\widehat{g}_{i\ell}^{(m)}.$ (58)
@@
$\max_{1\leq j_{1},j_{2}\leq r}\left|\frac{1}{T}\sum_{t=1}^{T}f_{j_{1},t}f_{j_{2},t}-\mathbb{E}f_{j_{1},t}f_{j_{2},t}\right|$ $\leq C_{1}\sqrt{\Theta_{j_{1},j_{1}}\Theta_{j_{2},j_{2}}}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right).$
@@
$\leq\frac{4\|A_{k}\|_{\rm 2}^{2}\max_{j\neq i}|\lambda_{j,i}|+4\|\Psi\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top}\|_{\rm 2}}{\alpha^{2k-2}\Theta_{ii}}.$ (63)
@@
$\Delta_{1,1}=\sum_{j_{2}\neq i}^{r}\frac{1}{T}\sum_{t=1}^{T}f_{it}f_{j_{2},t}\otimes_{\ell=1}^{K}a_{i\ell}\otimes_{\ell=K+1}^{2K}a_{j_{2}\ell}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top},$
$\Delta_{1,2}=\sum_{j_{1}\neq i}^{r}\frac{1}{T}\sum_{t=1}^{T}f_{j_{1},t}f_{it}\otimes_{\ell=1}^{K}a_{j_{1}\ell}\otimes_{\ell=K+1}^{2K}a_{i\ell}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top},$
$\Delta_{1,3}=\sum_{j_{1}\neq j_{2}\neq i}^{r}\frac{1}{T}\sum_{t=1}^{T}f_{j_{1},t}f_{j_{2},t}\otimes_{\ell=1}^{K}a_{j_{1}\ell}\otimes_{\ell=K+1}^{2K}a_{j_{2}\ell}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top}.$

### v2

@@ Appendix A Proofs of Main Theorem
$\overline{\mathbb{P}}\left(\left\|\sum_{t=1}^{T}\left[\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}-\overline{\mathbb{E}}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right]\right\|_{2}\geq x\right)$
$\leq$ $2d\exp\left(-\frac{x^{2}/2}{\sigma_{1}^{2}+N_{1}x/3}\right).$
Choosing $x\asymp\sqrt{d}\log(d)(\log T)^{(\vartheta+2)/\vartheta}\|\sum_{i=1}^{r}w_{i}f_{it}a_{i}\|_{2}+\sqrt{\log(d)}\sigma_{2}$, with probability at least $1-d^{-c_{4}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}-\overline{\mathbb{E}}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq$ $C_{4}\frac{\sqrt{\log(d)}}{T}\sigma_{2}+C_{4}\frac{\sqrt{d}\log(d)(\log T)^{1+\frac{2}{\vartheta}}\|\sum_{i=1}^{r}w_{i}f_{it}a_{i}\|_{2}}{T}.$
By Assumption 4.2, with probability at least $1-T^{-c_{5}}$,
$\left\|\sum_{i=1}^{r}w_{i}f_{it}a_{i}\right\|_{2}=\|F_{t}^{\top}WA^{\top}\|_{2}\leq(1+\delta)\|WF_{t}\|_{2}\lesssim\sqrt{r}(\log(T))^{1/\gamma_{1}}\sqrt{\lambda_{1}}.$
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq$ $C_{5}\sqrt{\frac{d\log(d)}{T}}\sqrt{\lambda_{1}}+C_{5}\cdot\frac{\sqrt{dr}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\lambda_{1}}}{T}.$
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\geq x\right)$
$\leq$ $\mathbb{P}(|M|>C)+\mathbb{P}\left(|M|<C,|M|\max_{t}\left\|\sum_{i=1}^{r}w_{i}f_{it}a_{i}\right\|_{2}\cdot\|e_{t}\|_{2}\geq x\right).$
Choosing $x\asymp\sqrt{dr}(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\lambda_{1}}$, we have with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}-T^{-c_{5}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\leq C_{6}\cdot\frac{\sqrt{dr}(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\lambda_{1}}}{T}.$
@@
$\|\Delta_{3}\|_{2}\leq$ $\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\leq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$+\left\|\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}+\left\|\mathbb{E}\sum_{i=1}^{r}w_{i}f_{i,t}a_{i}e_{t}^{\top}\mathbf{1}_{\{\|e_{t}\|_{2}\geq C\sqrt{d}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq$ $C_{7}\sqrt{\frac{d\log(d)}{T}}\sqrt{\lambda_{1}}+C_{7}\cdot\frac{\sqrt{dr}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\lambda_{1}}}{T}.$
@@
Let $\lambda_{1}/\lambda_{r}=\kappa$. Assume $d_{1}\lesssim\lambda_{1}$ and $\delta_{1}^{2}\kappa\leq c$ for a sufficiently small positive constant $c$. Apply random projection in Procedure 2 to the whole sample covariance tensor $\widehat{\Sigma}$ with $L\geq Cd^{2}\vee Cdr^{2\kappa^{2}}$. Denote the estimated CP basis vectors as $\widetilde{a}_{\ell k}$, for $1\leq\ell\leq L,1\leq k\leq K$. Then in an event with probability at least $1-T^{-c}-d^{-c}$, we have for any CP factor loading vectors tuple $(a_{ik},1\leq k\leq K)$, there exist $j_{i}\in[L]$ such that
@@
$\left\|\widetilde{a}_{j_{i},1}\widetilde{a}_{j_{i},1}^{\top}-a_{i1}a_{i1}^{\top}\right\|_{2}$ $\leq\psi_{i}+(\delta/\delta_{1})\kappa+\psi_{i}^{K-1}\kappa,$ (48)
where $\psi_{i}=CR^{(0)}/\Theta_{ii}+\delta_{1}^{2}\kappa$, $R^{(0)}=\phi^{(0)}$ is defined in (19) in Theorem 4.1, and $1\leq i\leq r$.
@@
Without loss of generality, assume $\Theta_{11}\geq\Theta_{22}\geq\cdots\geq\Theta_{rr}$. Then $\Theta_{rr}\geq\lambda_{r},\Theta_{11}\leq\lambda_{1}$. Let $\widehat{\Theta}_{ij}=T^{-1}\sum_{t=1}^{T}w_{i}w_{j}f_{it}f_{jt}$ and $\Sigma_{{\cal E}}=\mathbb{E}{\cal E}_{t}\otimes{\cal E}_{t}$. Write
@@
$\quad+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{it}\otimes_{k=1}^{K}a_{ik}\otimes{\cal E}_{t}+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{it}{\cal E}_{t}\otimes_{k=K+1}^{2K}a_{ik}+\left(\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes{\cal E}_{t}-\Sigma_{{\cal E}}\right)+\Sigma_{{\cal E}}$
@@
$\leq C\left(\sqrt{\frac{dd_{1}\log(d)}{T}}+\frac{d\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}\right)\cdot\sqrt{\log(d)},$
@@
$\left\|(\widetilde{\Delta}_{1}+\widetilde{\Delta}_{2}+\widetilde{\Delta}_{3}+\widetilde{\Delta}_{4})\times_{3}\times_{4}\theta\right\|_{2}$
$\leq$ $C\lambda_{1}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right)\cdot\sqrt{\log(d)}+C\left(\sqrt{\frac{dd_{1}\log(d)}{T}}+\frac{d\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}\right)\cdot\sqrt{\log(d)}$
$+C\lambda_{1}^{1/2}\left(\sqrt{\frac{d\log(d)}{T}}+\frac{\sqrt{dr}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}}{T}\right)\cdot\sqrt{\log(d)}.$
@@
$\geq(1-2\delta_{1}^{2}\kappa)\left(\sqrt{2\log(L)}-\frac{\log\log(L)}{4\sqrt{\log(L)}}-\sqrt{2\log(8)}\right)\Theta_{ii}-\left(\sqrt{4\log(r)}+\sqrt{2\log(8)}\right)\kappa\Theta_{ii}$
@@
with probability at least $\frac{1}{2}$, by letting $L\geq Cd\vee Cr^{2\kappa^{2}}$.
Since $\theta_{\ell}$ are independent samples, we instead take $L_{i}=L_{i1}+\cdots+L_{iM}$ for $M=\lceil C_{1}\log(d)/\log(2)\rceil$ and $L_{i1},...,L_{iM}\geq Cd\vee Cr^{2\kappa^{2}}$. We define
@@
$\left\|\widetilde{a}_{\ell_{*},k}\widetilde{a}_{\ell_{*},k}^{\top}-a_{ik}a_{ik}^{\top}\right\|_{2}\leq\frac{CR^{(0)}}{\Theta_{ii}}+\delta_{1}^{2}\kappa,\qquad 2\leq k\leq K.$ (53)
Now consider to obtain $\widetilde{a}_{\ell_{*},1}$. Write $\psi_{i}=CR^{(0)}/\Theta_{ii}+\delta_{1}^{2}\kappa$. Note that
@@
$\leq C_{K}(\delta/\delta_{1}+\psi_{i}^{K-1})\kappa\Theta_{ii}.$
@@
$\left\|\widetilde{a}_{\ell_{*},1}\widetilde{a}_{\ell_{*},1}^{\top}-a_{i1}a_{i1}^{\top}\right\|_{2}\leq\frac{CR^{(0)}}{\Theta_{ii}}+(\delta/\delta_{1})\kappa+\psi_{i}^{K-1}\kappa$ (55)
Repeat the same argument again for all $1\leq i\leq r$ factors, and let $L=\sum_{i}L_{i}\geq Cd^{2}\vee Cdr^{2\kappa^{2}}\geq Cdr\log(d)\vee Cr^{2\kappa^{2}+1}\log(d)$. We have, in the event $\Omega_{0}\cap\Omega_{1}\cap\cdots\cap\Omega_{r}$ with probability at least $1-T^{-c}-d^{-c}$, (53) and (55) hold for all $i$.
@@
Recall $\widehat{A}_{k}^{(m)}=(\widehat{a}_{1k}^{(m)},\ldots,\widehat{a}_{rk}^{(m)})\in\mathbb{R}^{d_{k}\times r}$, $\widehat{\Sigma}_{k}^{(m)}=\widehat{A}_{k}^{(m)\top}\widehat{A}_{k}^{(m)}$, and $\widehat{B}_{k}^{(m)}=\widehat{A}_{k}^{(m)}(\widehat{\Sigma}_{k}^{(m)})^{-1}=(\widehat{b}_{1k}^{(m)},...,\widehat{b}_{rk}^{(m)})\in\mathbb{R}^{d_{k}\times r}$. Let $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|F_{t},1\leq t\leq T)$ and $\overline{\mathbb{P}}(\cdot)=\mathbb{P}(\cdot|F_{t},1\leq t\leq T)$. Also let
$\overline{\lambda}_{i}=\frac{1}{T}\sum_{t=1}^{T}w_{i}^{2}f_{it}^{2}.$
@@
$=\sum_{i=1}^{r}\overline{\lambda}_{i}\otimes_{k=1}^{2K}a_{ik}+\sum_{i\neq j}^{r}\frac{1}{T}\sum_{t=1}^{T}w_{i}w_{j}f_{it}f_{jt}\otimes_{k=1}^{K}a_{ik}\otimes_{k=K+1}^{2K}a_{jk}+\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes{\cal E}_{t}$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{it}\otimes_{k=1}^{K}a_{ik}\otimes{\cal E}_{t}+\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{r}w_{i}f_{it}{\cal E}_{t}\otimes_{k=K+1}^{2K}a_{ik}$
@@
[…]
@@
[…]
@@
[…]
@@
$\max_{1\leq j_{1},j_{2}\leq r}\left|\frac{1}{T}\sum_{t=1}^{T}w_{j_{1}}w_{j_{2}}f_{j_{1},t}f_{j_{2},t}-\mathbb{E}w_{j_{1}}w_{j_{2}}f_{j_{1},t}f_{j_{2},t}\right|$ $\leq C_{1}\sqrt{\Theta_{j_{1},j_{1}}\Theta_{j_{2},j_{2}}}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right).$
@@
[…]
@@
$\Delta_{1,1}=\sum_{j_{2}\neq i}^{r}\frac{1}{T}\sum_{t=1}^{T}w_{i}w_{j_{2}}f_{it}f_{j_{2},t}\otimes_{\ell=1}^{K}a_{i\ell}\otimes_{\ell=K+1}^{2K}a_{j_{2}\ell}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top},$
$\Delta_{1,2}=\sum_{j_{1}\neq i}^{r}\frac{1}{T}\sum_{t=1}^{T}w_{i}w_{j_{1}}f_{j_{1},t}f_{it}\otimes_{\ell=1}^{K}a_{j_{1}\ell}\otimes_{\ell=K+1}^{2K}a_{i\ell}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top},$
$\Delta_{1,3}=\sum_{j_{1}\neq j_{2}\neq i}^{r}\frac{1}{T}\sum_{t=1}^{T}w_{j_{1}}w_{j_{2}}f_{j_{1},t}f_{j_{2},t}\otimes_{\ell=1}^{K}a_{j_{1}\ell}\otimes_{\ell=K+1}^{2K}a_{j_{2}\ell}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top}.$